integrate the following integral
∫ (x -1)/sqrt(x +4) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int \frac{x-1}{\sqrt{x+4}}dx
I = \int \frac{x+4-5}{\sqrt{x+4}}dx
I = \int [\sqrt{x+4}-\frac{5}{\sqrt{x+4}}]dx
I = \int \sqrt{x+4}dx-\int \frac{5}{\sqrt{x+4}}dx
I = \frac{2(x+4)^{3/2}}{3}-10\sqrt{x+4}+c